40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 29 16:25:06 2011, kiwamu, Update, ASC, beam axis and Y arm aligned 
    Reply  Sat Jul 30 06:24:06 2011, kiwamu, Update, ASC, Y arm ASS fixed 
       Reply  Sat Jul 30 21:04:23 2011, kiwamu, Update, ASC, Y arm ASS fixed 
       Reply  Sun Jul 31 00:05:52 2011, kiwamu, Update, LSC, script for loss measurement : modified 
          Reply  Sun Jul 31 22:35:35 2011, kiwamu, Update, LSC, arm loss measurement : done 
             Reply  Wed Sep 7 16:21:35 2011, kiwamu, Update, LSC, arm loss measurement : resluts Yarmloss.pngXarmloss.png
                Reply  Wed Sep 7 17:28:41 2011, kiwamu, Update, LSC, X arm loss measurement : not good Xarmlossmeasurement.png
Message ID: 5074     Entry time: Sun Jul 31 00:05:52 2011     In reply to: 5067     Reply to this: 5077
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: script for loss measurement : modified 

I modified the script armloss so that the channel names in the script are properly adopted to the new CDS.

Additionally I disabled the ETMX(Y)_tickle command in the script.

The tickle command puts some offsets on the LSC signal to let the arms pass through a fringe until it gets locked, but apparently we don't need it because the arms are loud enough.

A brief check showed that the script ran fine.

I will measure the loss on the X and Y arm cavity tomorrow.

Quote from #5067

Next : Health-check for the X arm ASS, the loss measurements.

 

ELOG V3.1.3-